It takes a boy 60 minutes to mow the lawn, but his sister can mow it in 90 minutes. How long would it take them to mow the lawn if they worked together, using two lawn mowers?
Answer by stanbon(75887)

It takes a boy 60 minutes to mow the lawn, but his sister can mow it in 90 minutes. How long would it take them to mow the lawn if they worked together, using two lawn mowers?
------------------
boy rate = 1/60 job/min
girl rate= 1/90 job/min
---
together rate = 1/x job/min
==========================
Equation:
rate + rate = together rate
1/60 + 1/90 = 1/x
Multiply thru by 180x to get:
3x + 2x = 180
5x = 180
x = 36 minutes (time to mow together)
==============
Cheers,
Stan H.